If a bullet traveling at 100m/s can penetrate 2 wooden block of same thickness.if it’s speed is doubled ,then what number of wooden blocks it can penetrate?

Grade:8

2 Answers

Arun
4 years ago
Dear student
 
 
3rd equation of motion can be be applied: v^2 = u^2 + 2as, where acceleration is -ve. Hence v^2 = u^2 – 2as.
v, final velocity = 0. Therefore, u^2 = 2as.
i.e. s is proportional to u^2. → s1÷ s2 = u1^2 ÷ u2^2
Therefore, s= (u2^2 ÷ u1^2) x s1 = (200÷100)^2 x 2 (planks) = 8 planks
Answer is 8
Khimraj
4 years ago

When speed is doubled, the kinetic energy is quadruple.

But…

In practice bullets never stay in the same shape when penetrating hard materials, like wood. They deform which alters the path and increases friction…

So theoretically I'd say 8 planks, practically I wouldn't know.